A father is 25 years older than his son. Today, the sum of their ages is 55 years. What is the age of the father and the son?
Answer by mananth(16949)

son's age be x
father's age = 25+x
sum of ages
x+x+25=55
2x+25=55
-25
2x+25-25=55-25
2x=30
/2
x=15
Son is 15
and father is 40 years